Two men start from two places 50 km apart.when they travel in opposite directions,they meet in two hours.when they travel in the same direction,one overtakes the other in five hours.Determine their respective rates of traveling
Answer by lwsshak3(11628)

Two men start from two places 50 km apart.when they travel in opposite directions,they meet in two hours.when they travel in the same direction,one overtakes the other in five hours. Determine their respective rates of traveling
***
let x=rate of faster man
let y=rate of slower man
rate=distance/travel time
..
x+y=rate travelling in opposite direction (toward each other)
x-y=rate travelling in same direction
..
2(x+y)=50
5(x-y)=50
..
2x+2y=50
5x-5y=50
..
10x+10y=250
10x-10y=100
add
20x=350
x=17.5
10y=250-10x=250-175=75
y=7.5
rate of faster man=17.5 mph
rate of slower man=7.5 mph